You have 125 mL of an 0.19 M acetic acid solution. What volume (V) of 1.10 M NaOH solution must you add in order to prepare an acetate buffer of pH = 4.20? (The pKa of acetic acid is 4.76.)
Answer #1

use the henderson hasselbalch equation:

pH=pKa+Log([A-]/[HA}])
[A-]/[HA] = 10^(4.22 - 4.76) = 0.2884
A- in your case is acetate
HA acetic acid

the number of moles of acetate and acetic acid can be used instead since the total volume is the same for both and concentration = moles/volume

number of moles of acetate = number of moles of NaOH which is [NaOH] x volume added (v) or [NaOH]v =1.1v

number of moles of acetic acid = initial moles added (0.125L x 0.19 M) - moles of NaOH (see above)
=0.02375-1.1v

pH=pKa+Log(1.1v/(0.2375-1.1v) )

1.1v/(0.2375-1.1v)=.288

1.1v=.0684-.3168V

1.4168v=.0684

v=.0483L=48.3ml